OpenStudy (anonymous):

t = 4π/3 (a) Find the reference number for the value of t.

OpenStudy (jdoe0001):

\(\bf t=\cfrac{5\pi}{3}\qquad \pi \approx 3.1416 \qquad \implies \cfrac{5\cdot 3.1416}{3}\)

OpenStudy (anonymous):

i tried that but it was wrong when i put it in

OpenStudy (jdoe0001):

sounds pretty ambiguous... can you post a screenshot of the material?

OpenStudy (jdoe0001):

hmm I used 5 it's 4, shoot a... anyhow

OpenStudy (jdoe0001):

do you have a Unit Circle?

OpenStudy (anonymous):

for this question?

OpenStudy (anonymous):

thats all there is to this question

OpenStudy (jdoe0001):

for any whatsoever.... you should, is handy for getting known angles well, many online http://mathref.weebly.com/uploads/7/6/5/6/7656473/3491886_orig.jpg

OpenStudy (jdoe0001):

notice where \(\bf \cfrac{4\pi}{3}\) is, the reference angle is the angle it'd make with the x-axis

OpenStudy (anonymous):

i know that. its just that every answer i put in is wrong

OpenStudy (anonymous):

wait nvm i got the answer. i just used the wrong pi symbol

OpenStudy (jdoe0001):

you use the angle the terminal point makes with the x-axis, and then "REFERENCE IT" to the 1st Quadrant

OpenStudy (anonymous):

thanks anyway

OpenStudy (jdoe0001):

yw